\begin{thebibliography}{8}

\bibitem{FG} S.Ju.Favorov and N.P.Girya,  {\it On a criterion of periodicity of polynomials,} Ufa's
 Mathematical Journal 4, No.1 (2012), 47-52.

\bibitem{FRR} S.Yu.\,Favorov, A.Yu.\,Rashkovskii, A.I.\,Ronkin {\it Almost periodic
divisors in a strip},  J. d'Analyse Math., Vol 74 (1998). P.
325--345.

\bibitem{G} N.P.Girya {\it Periodicity of Dirichlet series}, Matematychni studii, v.42,  ¹1, (2014), 13-18.

\bibitem{JT}  B.~Jessen, H.~Tornehave
 {\it Mean motions and zeros of almost periodic functions}
Acta Math. 77.  (1945), 137-279.

\bibitem{KL}
  M.G.Krein and B.Ja.Levin,
 {\it  On entire almost periodic functions
of an exponential type}, DAN SSSR v.64 (1949), No 2, 285--287 (in
Russian).

\bibitem{L} B.Ja.Levin,  {\it Distributions of Zeros of Entire Functions.}
Vol.5, Transl. of Math. Monograph,  AMS Providence, R1, 1980.

\bibitem{Lt}
B.M.Levitan  {\it Almost periodic functions}, Gostehizdat, Moscow,
1953 (in Russian).

\bibitem{T}
H.\,Tornehave, {\it Systems of zeros of holomorphic almost
periodic functions}, Kobenhavns Universitet Matematisk Institut,
Preprint no. 30, 1988.



\end{thebibliography}

\documentclass[12pt,a4paper]{amsart}

\usepackage {amsfonts}
\usepackage[english]{babel}
\def\g{\gamma}
\def\d{\delta}
\def\a{\alpha}
\def\b{\beta}
\def\e{\varepsilon}
\def\l{\lambda}
\def\R{{\mathbb R}}
\def\C{{\mathbb C}}
\def\N{{\mathbb N}}
\def\Z{{\mathbb Z}}
\def\Re{{\rm Re}}
\def\Im{{\rm Im}}
\def\dist{\mbox{dist}}

\def\supp{\mbox{supp}}

\newtheorem{theo}{Theorem}
\newtheorem{de}{Definition}

\newtheorem{Le}{Lemma}
\begin{document}

\title{Periodicity of almost periodic zero sets \\ and almost periodic functions}

\author{S.Yu. Favorov}

\address{Sergey Yuryevich Favorov,
\newline\hphantom{iii} The Karazin Kharkiv National University
\newline\hphantom{iii} Svoboda squ., 4,
\newline\hphantom{iii} 61022, Kharkiv, Ukraine}
\email{sfavorov@gmail.com}

\maketitle {\small
\begin{quote}
\noindent{\bf Abstract.} Whenever all differences between zeros of
two holomorphic almost periodic functions in a strip form a
discrete set, then both functions are infinite products of
periodic functions with commensurable periods. In particular, the
result is valid for some classes of Dirichlet series.
\medskip

AMS Mathematics Subject Classification: 30B50, 42A75

\medskip
\noindent{\bf Keywords:} quasipolynomial,  periodic function, zero
set, Dirichlet series, almost periodic holomorphic function.
\end{quote}
}

\bigskip
It was proved in \cite{FG} that each quasipolynomial
 \begin{equation}\label{Q}
   Q(z)=\sum_{n=1}^Na_ne^{\l_nz},\quad \l_n\in\R, \quad
   a_n\in\C.
 \end{equation}
with a discrete set of differences between its zeros is periodic
up to a multiplier without zeros, i.e., is of the form
 \begin{equation}\label{Q1}
Q(z)=Ce^{\b z}\prod_{k=1}^{N}\cosh(\omega
  z+b_k),\quad \b,\omega\in\R,
   \quad \ \C,\,
   b_k\in\C.
\end{equation}
The result is also valid for infinite sums
 \begin{equation}\label{S}
 S(z)=\sum_{n=1}^\infty a_ne^{\l_nz},\quad \l_n\in\R, \quad  a_n\in\C,
 \end{equation}
under conditions
 \begin{equation}\label{S1}
 \sum_{n=1}^\infty |a_n|<\infty, \quad \l_1=\sup_n\l_n<\infty,\quad \l_2=\inf_n\l_n>-\infty,\quad a_1a_2\neq0.
 \end{equation}
Note that zeros of  functions (\ref{Q}), (\ref{S}) are located
in a vertical strip of a finite width.

In 1949 M.G. Krein and B.Ja.Levin (see \cite{KL}, also \cite{L},
 ch.6, p.2 and Appendix~6) introduced and studied the class
$\Delta$ of entire almost periodic functions of an exponential
growth with zeros in a horizontal strip of a finite width. In
particular, if a sum $S(z)$ in (\ref{S}) satisfies condition
(\ref{S1}), then $S(iz)$ belongs to $\Delta$. In fact,
representation (\ref{Q1}) was obtained in \cite{FG} for functions
from $\Delta$ with a discrete sets of differences between zeros
(of course, it should be replaced $e^{\b z}$ by $e^{i\b z}$ and
$\cosh(\omega z+b_k)$ by $\cos(\omega z+b_k)$ ).

The phenomenon takes place for a pair of functions and discreteness of the set of differences between their zeros as well
(\cite{G}). For example, let $Q_1,\,Q_2$ be entire functions of the form (\ref{S}) under conditions (\ref{S1}).
 If the set $\{z-w,\ Q_1(z)=0,\,Q_2(w)=0\}$ is discrete, then the both functions have
the form (\ref{Q1}) with the same $\omega$ and possibly different $C,\b,N,b_k$. In particular, in the case $Q_2(-z)=Q_1(z)=Q(z)$ we
get representation (\ref{Q1}) for any function (\ref{S}) with a discrete set $\{z+z',\ Q(z)=Q(z')=0\}$.

The proof of the above results are based on the  property of zeros
$\{z_j\}$ of functions $f\in\Delta$ (\cite{L}, Appendix~6, p.2) to
be almost periodic in the  sense of the following definition:
\begin{de}
A zero set $\{z_j\}$ is almost periodic if for any $\e>0$ there is a
relatively dense set $E_\e\subset\R$  such that for each $\tau\in
E_\e$ there exists a bijection $\sigma:\,\N\to\N$ with the
property
 \begin{equation}\label{ap2}
\sup_j |z_j+i\tau-z_{\sigma(j)}|<\e.
 \end{equation}
 \end{de}
 Recall that a set
$E\subset\R$ is relatively dense, if there exists $L<\infty$ such
that $E\cap[a,a+L]\neq\emptyset$ for any $a\in\R$.

\medskip
Practically, the above definition is applied to zero sets in a {\it closed} vertical strip.
In the present paper we investigate holomorphic
almost periodic functions and almost periodic sets in an open
strip, in particular, in the complex plane and in half-planes.

\begin{de}
A continuous function $f(z)$ in a strip $S=\{z:\,a<\Re z<b\},\
-\infty\le a<b\le\infty$ is almost periodic, if for any substrip
$S_0$, $\overline{S_0}\subset S$ and any $\e>0$ there is a
relatively dense set $E_{\e,S^0}\subset\R$ such that for any
$\tau\in E_{\e,S^0}$
 \begin{equation}\label{apf}
 \sup_{z\in S_0}|f(z+i\tau)-f(z)|<\e.
 \end{equation}
\end{de}
A typical example of an entire almost periodic function is sum
(\ref{S}) under  conditions
 $$
 \sum_{n=1}^\infty |a_n|<\infty,\quad \sup_n\l_n<\infty,\quad \inf_n\l_n>-\infty.
 $$
  Moreover, a holomorphic function $f(z)$ in a strip
$S=\{z:\,a<\Re z<b\},\ -\infty\le a<b\le\infty$ is almost periodic
if and only if there exists a sequence of quasipolynomials $Q_n$
of the form (\ref{Q}) such that  for any substrip $S_0$,
$\overline{S_0}\subset S$,
 $$
 \sup_{z\in S_0}|f(z)-Q_n(z)|\to0,\quad n\to\infty.
 $$
(see \cite{JT}, item 8).

In order to take into account multiplicities of zeros, we  use
the term {\it divisor} instead of zero set. Namely, a divisor $Z$ in
a domain $D$ is a mapping $Z:\,D\to\N\cup\{0\}$ such that
$|Z|=\supp Z$ is a set without limit points in $D$. In other
words, a divisor in $D$ is the sequence of points $\{z_j\}\subset
D$ that has no limit points in $D$ and every point of $D$ appears at
most a finite times in the sequence. If $Z(z)\le1$ for all $z\in D$,
we will identify $Z$ and $|Z|$. The divisor of zeros of a
holomorphic function $f$ in $D$ is the map $Z_f$ such that
$Z_f(a)$ equals multiplicity of zero of function $f$ at the point
$a$.

The following definition appeared at first in \cite{T}:

\begin{de}[\cite{T},\cite{FRR}]\label{d3} A divisor $Z=\{z_j\}$ in a strip
$S=\{z:\,a<\Re z<b\},\ -\infty\le a<b\le\infty$, is called almost
periodic if for any $\e>0$ and any substrip $S^0$,
$\overline{S^0}\subset S$, there is a relatively
dense set
 \begin{equation}\label{ap3}
 E_{\e,S^0}=\{\tau\in\R:\,z_j\in S^0 \bigvee z_{\sigma(j)}\in S^0
\Longrightarrow |z_j+i\tau-z_{\sigma(j)}|<\e\}.
 \end{equation}
where $\sigma=\sigma_\tau$ is a suitable bijection $\N\to\N$.
\end{de}

{\bf Remark.} If we take $\sigma^{-1}(j)$ instead of $j$ in
(\ref{ap3}), we get
\begin{equation}\label{ap1}
z_j\in S^0 \bigvee z_{\sigma^{-1}(j)}\in S^0 \Longrightarrow
|z_j-i\tau-z_{\sigma^{-1}(j)}|<\e.
 \end{equation}

\begin{theo}[\cite{FRR}]\label{t1}
a) If $f$ is a holomorphic function in a strip $S$ with an almost
periodic modulus, then the divisor of zeros of $f$ is almost
periodic,

b) for any almost periodic divisor $Z$ in a strip $S$ there is a
holomorphic function $f$ in $S$ with an almost periodic modulus
such that $Z=Z_f$,

c) for any almost periodic divisor $Z$ in a strip $S$ such that
$|Z|\cap S'=\emptyset$ for some open substrip $S'\subset S$ there
is a holomorphic almost periodic function $f$ in $S$ such that
$Z=Z_f$.
\end{theo}

 \medskip
 There exist almost periodic divisors in the plane with a
 discrete set of differences that is not periodic.

 {\bf Example 1.} Let
 $$
 Z=\{z_{n,k}=i2^{nk}+2^k,\ n\in\Z,\ k\in\N\}
 $$
be a discrete set in $\C$. It is easy to see that the set is
almost periodic, differences between its points form a discrete
set, but $Z$ is a countable union of periodic sets with different
commensurable periods.

\medskip
In our article we prove the following theorem

\begin{theo}\label{t2}
Let $Z$, $W$ be  almost periodic divisors in a strip
$S=\{z:\,a<\Re z<b\},\ -\infty\le a<b\le\infty$. If

a) $|Z|\cap S'=\emptyset$ for some substrip $S'\subset S$,

b)  for any substrip $S_0$ of a finite width,
$\overline{S_0}\subset S$, the set $\{z-w,\ z\in |Z|\cap
S_0,\,w\in |W|\cap S_0\}$ is discrete,

\noindent then $Z,\,W$ are at most countable sums of periodic
divisors with commensurable periods. \end{theo}

Show that condition a) is essential.

{\bf Example 2.} Let $S=\{z:\,|\Re z|<1\},\
Z=\{z_{m,n}=(m+in)e^{i\a}\in S,\,m,n\in\Z\}$, where $\a\in\R$ such
that $\cot\,\a$ is an irrational number. Clearly, the set of
differences of elements of $Z$ is discrete. Let us prove that $Z$
is an almost periodic divisor without any periods.

By Kronecker Lemma (see, for example, \cite{Lt}, Ch.2, \$ 2), for
any $\e>0$ the inequalities
$$
|t\cot\a|<\e\ (\text{mod}\,\Z),\quad |t|<\e\ (\text{mod}\,\Z)
 $$
 has a relatively dense set of
common solutions. Therefore, the inequality
$$
|m\cot\a|<\e(1+|\cot\a|)\ (\text{mod}\,\Z)
$$
has a relatively dense set of integer solutions. The later means
that for any $\d$ there exist pairs of integers $(m,n)\in\Z^2$
such that
 \begin{equation}\label{in}
\quad |m\cos\a-n\sin\a|<\d,
\end{equation}
and the set of $m$ with this property is relatively dense. Put
 $$
E=\{m\sin\a+n\cos\a:\,|m\cos\a-n\sin\a|<\d\}.
 $$
Let $\tau=\tilde m\sin\a+\tilde n\cos\a\in E$. For any $z_{m,n}\in
Z$ and $m'=m+\tilde m,\ n'=n+\tilde n$ we have
 $$
z_{m,n}+i\tau-z_{m',n'}=(m+in)e^{i\a}+i\Im[(\tilde m+i\tilde
n)e^{i\a}]-(m'+in')e^{i\a}=\Re[(\tilde m+i\tilde n)e^{i\a}].
 $$
By (\ref{in}), $|z_{m,n}+i\tau-z_{m',n'}|<\d$. Since
$$
|\tau-(\cos^2\a/\sin\a+\sin\a)\tilde m|=|\tilde n\cos\a-(\cos^2\a/\sin\a)\tilde m| <\d|\cot\a|,
$$
 we see that the set $E$ is
relatively dense. So, $Z$ is an almost periodic divisor.

 Since $Z$ sites in the vertical strip of width $2$, we see
that any period of $Z$ must have the form $iT,\ T\in\R$. Hence,
for some $(n,m),\,(n',m')\in\Z^2,\ (n,m)\neq(n',m')$,
 $$
(n+im)e^{i\a}-(n'+im')e^{i\a}=iT,
 $$
and
$$
(n-n')+i(m-m')=iT(\cos\,\a-i\sin\,\a).
 $$
This equality contradicts our choice of $\a$. Hence any part of
$Z$ has no periods.

\medskip
Our proof of Theorem \ref{t2} makes into use the following Lemmas

\begin{Le}\label{L1}
Let $Z=\{z_n\}$ be an almost periodic divisor in a strip $S$,
$S^0$ be a substrip, $\overline{S_0}\subset S,\ \e>0,\
\tau_1,\,\tau_2\in E_{\e,S^0}$, where $E_{\e,S^0}$ is from Definition \ref{d3}.
 Then there is a bijection $\sigma:\N\to\N$ such that
if $z_j\in S^0,\,\dist(z_j,\partial S^0)>\e$, then
\begin{equation}\label{ap0}
 |z_j+i(\tau_1-\tau_2)-z_{\sigma(j)}|<2\e,\quad
 |z_j-i(\tau_1-\tau_2)-z_{\sigma^{-1}(j)}|<2\e.
\end{equation}
 \end{Le}
It can be proved easily that the similar assertion is valid for
$\tau_1+\tau_2$ as well.

{\bf Proof.} If $z_j$ satisfies conditions of the Lemma, then for
some bijection $\sigma_1:\N\to\N$ we have
$|z_j+i\tau_1-z_{\sigma_1(j)}|<\e$, therefore
$|\Re(z_{\sigma_1(j)}-z_j)|<\e$ and $z_{\sigma_1(j)}\in S^0$. By
(\ref{ap1}), we also get the inequality
$|z_{\sigma_2^{-1}\circ\sigma_1(j)}-i\tau_2-z_{\sigma_1(j)}|<\e$ for some bijection $\sigma_2:\N\to\N$.
Therefore, we obtain the first part in (\ref{ap0}) with
$\sigma=\sigma_2^{-1}\circ\sigma_1$. If we change places of $\tau_1$ and
$\tau_2$, we obtain the second part in (\ref{ap0}).

 \begin{Le}\label{L2} Let $Z=\{z_n\},\,W=\{w_m\}$ be almost periodic
divisors in a strip $S$.  Then for any $\e>0$ and substrip
$S^0\subset S$, $\dist(S_0,\partial S)>\e$, there is a relatively
dense set $E\subset\R$ with the  property:  for any $\tau\in E$
there are bijections $\sigma_Z,\,\sigma_W$ from $\N$ to $\N$ such
that for all $z_j\in |Z|\cap S^0,\,w_r\in |W|\cap S^0$
\begin{equation}\label{ap4}
 |z_j+i\tau-z_{\sigma_Z(j)}|<\e,\quad
 |w_r+i\tau-w_{\sigma_W(r)}|<\e
 \end{equation}
 and
\begin{equation}\label{ap5}
 |z_j-i\tau-z_{\sigma^{-1}_Z(j)}|<\e,\quad
 |w_r-i\tau-w_{\sigma^{-1}_W(r)}|<\e.
 \end{equation}
 \end{Le}

{\bf Proof.} Put $\tilde S=\{z\in S:\,\dist(z,S^0)<\e/2\}$. By
(\ref{ap3}) and (\ref{ap1}), there is a number $L$ such that any
interval of length $L$ contains $\tau_Z$ and $\tau_W$ with the
properties
 \begin{equation}\label{ap6}
 |z_j+i\tau_Z-z_{\sigma_Z(j)}|<\e/4,\quad
 |z_j-i\tau_Z-z_{\sigma^{-1}_Z(j)}|<\e/4
 \end{equation}
 for all $z_j\in |Z|\cap\tilde S$, and
 \begin{equation}\label{ap7}
 |w_r+i\tau_W-w_{\sigma_W(r)}|<\e/4,\quad \quad
 |w_r-i\tau_W-w_{\sigma^{-1}_W(r)}|<\e/4
 \end{equation}
 for all $w_r\in |W|\cap\tilde S$. Here $\sigma_Z,\,\sigma_W$ are some
 (in general, different) bijections $\N\to\N$.

 We may suppose that $N=2L/\e$ is an integer. Hence for
any $k\in\Z$ there are integers $n(k),\,m(k),\,0\le n(k),m(k)\le
N$ such that
 $$
 |kL+n(k)\e/2-\tau_Z|<\e/4,\quad |kL+m(k)\e/2-\tau_W|<\e/4.
 $$
  The differences $n(k)-m(k)$ take at most $2N+1$ values. Choose
$k_1,\dots,k_r,\,r\le 2N+1,$ such that for any $k\in\Z$ there is
$k_s$ such that
 $$
 n(k)-m(k)=n(k_s)-m(k_s).
 $$
It follows easily that any interval of length $L(\max_s(|k_s|+2)$
contains a point of the form
 $$
 \tau=Lk+n(k)\e/2-Lk_s-n(k_s)\e/2=Lk+m(k)\e/2-Lk_s-m(k_s)\e/2.
 $$
If we replace $\tau_Z$ by $Lk+n(k)\e/2$ or $Lk_s+n(k_s)\e/2$ in
(\ref{ap6}), we obtain that this  inequality  satisfies with $\e/2$ instead
of $\e/4$.  The same is true if we replace $\tau_W$ by
$Lk+m(k)\e/2$ or $Lk_s+m(k_s)\e/2$ in (\ref{ap7}). Applying Lemma
\ref{L1} with $S^0=\tilde S$ and $\e/2$ instead of $\e$, we see
that $\tau$ satisfies (\ref{ap4}) and (\ref{ap5}).

 {\bf Proof of Theorem \ref{t2}.} Set a sequence of vertical substrips
$S_k,\,\overline{S_k}\subset S_{k+1},$ such that
 $$
|Z|\cap S_1\neq\emptyset,\quad |W|\cap S_1\neq\emptyset,\quad
S_1\cap S'\neq\emptyset,\quad\cup_k S_k=S.
 $$
 Take any points $z\in |Z|\cap S_1$, $w\in |W|\cap S_1$. It follows from
(\ref{ap4}) with $\e<\min\{\dist(z,\,\partial
S_1),\dist(w,\,\partial S_1)\}$ that there is $R<\infty$ such that
any horizontal strip of the width $R$ contains at least one point
$|Z|\cap S_1$ and at least one point $|W|\cap S_1$. By condition
b) of the theorem,  for each $k$ the set $\{z-w,\ z\in |Z|\cap
S_k,\,w\in |W|\cap S_k\}$ is discrete.  Hence there exists
$\g_k\in(0,\,1/2)$ such that whenever
$$
z_n,z_{n'}\in |Z|\cap S_{k+1},\ w_m,w_{m'}\in |Z|\cap S_{k+1},\quad z_n-w_m\neq z_{n'}-w_{m'},
$$
$$
|\Im(z_n-w_m)|<2R+3,\quad|\Im(z_{n'}-w_{m'})|<2R+3,
$$
 we get $\g_k<|(z_{n'}-w_{m'})-(z_n-w_m)|$. In particular, if we
put $w_m=w_{m'}$, then we get $\g_k<|z_n-z_{n'}|$ for any
$z_n,z_{n'}\in |Z|\cap S_{k+1}$, $z_n\neq z_{n'}$.

Fix $z_n\in |Z|\cap S_k$, and let a  number $\tau>1$ satisfies
(\ref{ap4}) for $Z$ and $W$ with $\e=\g_k/2$. Then there is a
unique
 $z_{n'}\in |Z|$ such that $|z_n+i\tau-z_{n'}|<\g_k/2$.
Indeed, otherwise we obtain
 $$
 |z_{n'}-z_{n''}|\le |z_n+i\tau-z_{n'}|+|z_n+i\tau-z_{n''}|<\g_k.
 $$
Set $T_k=(z_{n'}-z_n)/i$.

 Let $w_m\in |W|$ be such that
$|\Im\,(w_m-z_n)|<2R+2$. By (\ref{ap4}), there is a point
$w_{m'}\in |W|$ such that $|w_m+i\tau-w_{m'}|<\g_k/2$. Therefore,
  $$
|(z_n-w_m)-(z_{n'}-w_{m'})|\le|w_m+i\tau-w_{m'}|+|z_{n'}-z_n-i\tau|<\g_k.
  $$
   Since
  $$
|\Im(z_{n'}-w_{m'})|\le|\Im(z_n-w_m)|+|z_n-z_{n'}+i\tau|+|w_m-w_{m'}+i\tau|<2R+3,
 $$
  we get $z_n-w_m=z_{n'}-w_{m'}$ due to the choice of $\g_k$.
 Therefore, $w_{m'}=w_m+iT_k$.

 The latter equality takes place for all points
$|W|\cap\{w:\,\Im z_n-2R<\Im w<\Im z_n+2R\}$, in particular, for
some $w_l$ such that $\Im z_n+R<\Im w_l<\Im z_n+2R$. Namely, there
is $w_{l'}\in |W|$ such that $w_{l'}=w_l+iT_k$. Let $\zeta\in |Z|$
be any point from the set
\begin{equation}\label{set}
\{z:\,\Im z_n\le\Im z<\Im z_n+3R\}\subset\{z:\,\Im w_l-2R<\Im
z<\Im w_l+2R\}.
\end{equation}

 By (\ref{ap4}), there is a point
$\zeta'\in |Z|$ such that $|\zeta+i\tau-\zeta'|<\g_k/2$.
Therefore,
  $$
|(\zeta-w_l)-(\zeta'-w_{l'})|\le|\zeta+i\tau-\zeta'|+|iT_k-i\tau|<\g_k.
  $$
   Since $|\Im\zeta-\Im w_l|<2R$ and
  $$
|\Im(\zeta'-w_{l'})|\le|\Im(\zeta-w_l)|+|\zeta+i\tau-\zeta'|+|iT_k-i\tau|<2R+3,
 $$
  we get $\zeta-w_l=\zeta'-w_{l'}$ due to the choice of $\g_k$.
 Therefore, $\zeta'=\zeta+iT_k$.

In particular, there is a point $z_s\in |Z|\cap\{z:\,\Im
z_n+2R\le\Im z<\Im z_n+3R\}$ such that $z_s+iT_k\in |Z|$.
Continuing the line of reasoning, we obtain that for all $z\in
|Z|$ such that $\Im z\ge\Im z_n$ we get $z+iT_k\in |Z|$ and for
all $w\in |W|$ such that $\Im w\ge\Im z_n$ we get $w+iT_k\in |W|$.

If we take $w'_l\in |W|$ such that $\Im z_n-2R<\Im w'_l<\Im
z_n-R$, we also can find $w'_{l'}\in |W|$ such that
$w'_{l'}=w'_l+iT_k$. Next, we prove that for any point $\tilde
\zeta\in |Z|\cap\{z:\,\Im z_n-3R\le\Im z<\Im z_n\}$ there is
$\tilde\zeta'\in |Z|$ such that $\tilde\zeta'=\tilde\zeta+iT_k$.
Arguing as above, we show that for all $z\in |Z|$ such that $\Im
z\le\Im z_n$ we get $z+iT_k\in |Z|$ and for all $w\in |W|$ such
that $\Im w\le\Im z_n$ we get $w+iT_k\in |W|$.

Next, by (\ref{ap5}), take for any $z\in |Z|$ a point  $z''\in
|Z|$ such that $|z''+i\tau-z|<\g_k/2$. Then $z''+iT_k\in |Z|$ and
 $$
|(z''+iT_k)-z|\le |z''+i\tau-z|+|i\tau-iT_k|<\g_k.
 $$
Therefore, $z''+iT_k=z$ and $z-iT_k\in |Z|$ for all $z\in |Z|$. By
the same arguments, $w-iT_k\in |W|$ for all $w\in |W|$.

If an imaginary part of $T_k$ does not vanish, then  either
$z_n+iMT_k\in S'$, or $z_n-iMT_k\in S'$ for a suitable $M\in\N$. Therefore, $T_k$ is real and for any $M\in\Z$ we get
$(|Z|\cap S_k)+iMT_k=|Z|\cap S_k,\ (|W|\cap S_k)+iMT_k=|W|\cap
S_k$. Hence, the restrictions $Z\mid_{S_k},\,W\mid_{S_k}$ the
divisors $Z,\,W$ to $S_k$ are periodic divisors with period
$iT_k$.

The same arguments work for every $k=1,2,\dots$.

Let $T^0_k$ be the minimal common period of $Z\mid_{S_k}$ and
$W\mid_{S_k}$. Clearly, $T_k/T^0_k\in\N$. Besides, since $|Z|\cap
S_k\subset |Z|\cap S_m$ for $m>k$, we have $T_m/T^0_k\in\N$ as
well.

Finally, let $Z_k=Z\mid_{S_{k+1}\setminus S_k},\
W_k=W\mid_{S_{k+1}\setminus S_k}$. Then we obtain
 $$
 Z=Z_1+Z_2+Z_3+\dots,\qquad W=W_1+W_2+W_3+\dots.
 $$
Theorem is proved.

Theorem \ref{t2} implies the corresponding result for almost
periodic holomorphic functions.
\begin{theo}\label{t3}
Let $f$, $g$ be  almost periodic functions in a strip
$S=\{z:\,a<\Re z<b\}, -\infty\le a<b\le\infty$. If

a) either $f$, or $g$ has no zeros in an open substrip $S'\subset
S$,

b)  for any substrip $S_0$, $\overline{S_0}\subset S$, the set
$\{z-w,\ z,w\in S_0,\,f(z)=g(w)=0\}$ is discrete, then
\begin{equation}\label{pro}
f(z)=f_0(z)\prod_{k=1}^\infty f_k(z),\quad
g(z)=g_0(z)\prod_{k=1}^\infty g_k(z)
\end{equation}
where  $f_0,\,g_0$ are holomorphic almost periodic functions in
$S$ without zeros,  $f_k,\,g_k,\ k=1,2,\dots,$
 are periodic holomorphic in $S$ with commensurable periods $T_k$.
\end{theo}

{\bf Proof.} By Theorem \ref{t1}a), the divisors $Z,\,W$ of zeros of $f,\,g$, respectively, 
 are almost periodic and satisfy other conditions of the previous theorem. Let
$S_k,\,Z_k,\,W_k$, be the same as
in the proof of Theorem \ref{t2}. Set
 $$
 S_k=\{z:\,\eta_k<\Re\,z<\eta'_k\},\quad\eta_k<\eta_{k-1},\quad\eta'_k>\eta'_{k-1},\ \forall k.
 $$
For any $k$ there is only a finite number of points
$a^k_1,\dots,a^k_{m_k}\in |Z_k|\cup\{z:\,0\le\Im z
<T_k\}$.  Therefore, $|Z_k|=\{a_1^k,\dots,a_{m_k}^k\}+iT_k\Z$. Take
$\e_k^j>0,\,1\le j\le m_k,\,3\le k<\infty,$ such that
 \begin{equation}\label{s}
\sum_{j,k}\e_k^j<\infty.
 \end{equation}
Fix $k>1$. Let
 $$
 h_j(w)=1-w\exp(-2\pi a^k_j /T_k ),\quad
j=1,\dots,m_k.
$$
Clearly, $h_j(\exp(2\pi z/T_k))$ has a divisor $a_j^k+iT_k\Z$.
Since $\a_j^k\in S_k\setminus S_{k-1`}$, we see that either $\eta'_{k-1}\le\Re\,a_j^k$, or
$\eta_{k-1}\ge\Re\,a_j^k$. In the first case,  $\log h_j(w)$ is
holomorphic on the disc $|w|<\exp(2\pi\Re\,a^k_j /T_k)$, and there
is a polynomial $P_j^k$  such that
 $$
|\log h_j(w)-P_j^k(w)|<\e_k^j\quad\text{for}\  |w|\le\exp(2\pi
\eta'_{k-2}/T_k).
 $$
In the second one, $\log h_j(w)$ is holomorphic on the set
$|w|>\exp(2\pi\Re\,a^k_j /T_k)$, and there is a polynomial $P_j^k$
 such that
 $$
|\log h_j(w)-P_j^k(w^{-1})|<\e_k^j\quad\text{for}\
|w|\ge\exp(2\pi \eta_{k-2}/T_k).
 $$
Put $Q_j(z)=P_j^k(\exp(2\pi z/T_k))$ in the first case, and
$Q_j(z)=P_j^k(\exp(-2\pi z/T_k))$ in the second one. We obtain
 \begin{equation}\label{apr}
|h_j(e^{2\pi z/T_k})e^{-Q_j(z)}|<e^{\e_j^k}\quad\text{for}\  z\in
S_{k-2}.
 \end{equation}
Put
 $$
 f_k(z)=\prod_{j=1}^{m_k}h_j(e^{2\pi z/T_k}),\ k=1,2,\quad
 f_k(z)=\prod_{j=1}^{m_k}h_j(e^{2\pi z/T_k})e^{-Q_j(z)},\ k>2.
 $$
Clearly, $Z_{f_k}=Z_k$. By (\ref{s}) and (\ref{apr}), the product
in (\ref{pro}) converges in every substrip $S_k$. The function
$f(z)/\prod_{k=1}^\infty f_k(z)$ is holomorphic in $S$, and, by
\cite{T}, Theorem 1, it is almost periodic in $S$. In the same way,
we obtain the representation of $g$.

{\bf Remark.} It follows easily that for entire almost periodic
functions $f,\,g$ with zeros in a strip $\tilde S$ of a finite
width one can take $S_1=\tilde S$. Therefore, $f,\,g$ are periodic
functions with the same period up to almost periodic multiplies
without zeros. Hence the results of \cite{FG}, \cite{G} follow
from Theorem \ref{t3}.

\end{document}